OpenStudy (anonymous):

Choose the equation below that represents the line passing through the point (-3, -1) with a slope of 4.

jimthompson5910 (jim_thompson5910):

y = mx+b y = 4x + b ... Plug in the given slope, m = 4 -1 = 4(-3) + b ... Plug in x = -3 and y = -1 from the given point (-3,-1) Now solve for b

jimthompson5910 (jim_thompson5910):

Tell me what you get

OpenStudy (anonymous):

y=4x-11

jimthompson5910 (jim_thompson5910):

close, but not quite

jimthompson5910 (jim_thompson5910):

y = mx+b y = 4x + b ... Plug in the given slope, m = 4 -1 = 4(-3) + b ... Plug in x = -3 and y = -1 from the given point (-3,-1) -1 = -12 + b -1+12= b 11 = b b = 11 The 11 is positive since -1+12 is the same as 12-1

OpenStudy (anonymous):

y=4x+11

jimthompson5910 (jim_thompson5910):

you got it

OpenStudy (anonymous):

Thank You So Much But I Have An Couple More That I Need Help On If You Don't Mind

jimthompson5910 (jim_thompson5910):

alright

OpenStudy (anonymous):

What is Correlation?

jimthompson5910 (jim_thompson5910):

it's like a connection between two or more variables

jimthompson5910 (jim_thompson5910):

but it's not a strong connection like one causes the other ex: there is a correlation between studying and getting good grades. As one goes up, the other goes up (but the two aren't directly connected)

OpenStudy (anonymous):

Ex House Address vs test score

jimthompson5910 (jim_thompson5910):

if you increase your house address, will that change your test score? How about if you decrease your house address?

OpenStudy (anonymous):

Okay AGE VS COMMUTE TIME TO SCHOOL

jimthompson5910 (jim_thompson5910):

same question: if you change your age, will that change the commute time?

OpenStudy (anonymous):

but as a positive correlation

jimthompson5910 (jim_thompson5910):

not sure what you mean

OpenStudy (anonymous):

That Alright Last Question How Would You Writethe equation of the line in slope-intercept form with a slope of -5 and a y-intercept of -1.

jimthompson5910 (jim_thompson5910):

y = mx+b is the general equation of a line

jimthompson5910 (jim_thompson5910):

where m is the slope and b is the y-intercept

OpenStudy (anonymous):

That's All

jimthompson5910 (jim_thompson5910):

the slope is -5 so m = -5 what is b?

OpenStudy (anonymous):

-1

jimthompson5910 (jim_thompson5910):

so y = mx+b becomes y = -5x - 1

OpenStudy (anonymous):

Okay Thank You Hopefully When I Need Help You Can Help Me

jimthompson5910 (jim_thompson5910):

yw